owenjiang:

Every female bear has 3 baby cubs. What is the constant of proportionality for the ratio of cubs to female bears?

owenjiang:

help pls

TWALKER16:

are there any question answers to this?

TWALKER16:

Is this a test

owenjiang:

1/2, 1/3,3,2

owenjiang:

answers

owenjiang:

it's 3 then 2

TWALKER16:

@Elsa213

jimthompson5910:

Let's say every female had 4 cubs This means we have the ratio 1 female: 4 cubs The constant of proportionality is the ratio of the number of cubs over the number of females So we say 4/1 = 4 is the constant of proportionality The equation is y = 4x x = number of females y = number of cubs eg: x = 2 female bears means y = 4x = 4*2 = 8 cubs total

sammixboo:

Everyone, please provide the steps and guide the user rather than give out the answers. Also, do not comment unless you are contributing to the original post. Thank you.

Jackalope:

what 2 numbers are in the equatio then thats ur answer